OpenStudy (anonymous):

please help: solve and check the equation: x/8=(x/9)-11

OpenStudy (anonymous):

x x -- = --- - 11 8 9 x x --- - --- = - 11 8 9 9x - 8x -------- = - 11 72 x ---- = -11 72 x = -11 * 72 x = - 792